OpenStudy (anonymous):

dy/dx= 3x^3y Solve the differential equation. Write the answer as an explicit solution (i.e. solve for the dependent variable).

OpenStudy (anonymous):

it is \[3x ^{3}*y\]

OpenStudy (anonymous):

u sure abt the power 3x^3y..i dont think i hv seen something like that before

Parth (parthkohli):

\[{dy \over dx} = 3x^{\large3}y\]Actually, that's the question for everyone's clarification.

OpenStudy (anonymous):

ans is ln(y)= 3x^4/4 +c

OpenStudy (anonymous):

ryt

OpenStudy (anonymous):

\[\frac{ dy }{ y} = { x ^{3}dx}\]

OpenStudy (anonymous):

yes integrate that and u get what i posted

OpenStudy (anonymous):

give me a medal psk if the ans is ryt :)'

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

r u moderator

OpenStudy (anonymous):

no u?

OpenStudy (anonymous):

\[\frac{dy}{dx} =3x^{3} y => \frac{dy}{y} =3x^3dx => \int\limits\frac{dy}{y} =\int\limits3x^{3}dx => \ln(y)=\frac{3x^{4}}{4} +C\] i think you have to find the analytical solution next

OpenStudy (anonymous):

dy/dx= 3x^(3)*y 3.3x2. how about y??

OpenStudy (anonymous):

then if i take ln of both sides i should get the whole right side in exponent form??